# RECON: Reducing Causal Confusion with Human-Placed Markers

### v1

@@ Abstract
Imitation learning enables robots to learn new tasks from human examples. One current fundamental limitation while learning from humans is causal confusion. Causal confusion occurs when the robot’s observations include both task-relevant and extraneous information: for instance, a robot’s camera might see not only the intended goal, but also clutter and changes in lighting within its environment. Because the robot does not know which aspects of its observations are important a priori, it often misinterprets the human’s examples and fails to learn the desired task. To address this issue, we highlight that — while the robot learner may not know what to focus on — the human teacher does. In this paper we propose that the human proactively marks key parts of their task with small, lightweight beacons. Under our framework the human attaches these beacons to task-relevant objects before providing demonstrations: as the human shows examples of the task, beacons track the position of marked objects. We then harness this offline beacon data to train a task-relevant state embedding. Specifically, we embed the robot’s observations to a latent state that is correlated with the measured beacon readings: in practice, this causes the robot to autonomously filter out extraneous observations and make decisions based on features learned from the beacon data. Our simulations and a real robot experiment suggest that this framework for human-placed beacons mitigates causal confusion and enables robots to learn the desired task from fewer demonstrations. See videos here: https://youtu.be/oy85xJvtLSU
@@ I Introduction
Leveraging Beacons. We extend imitation learning from offline human demonstrations to include user-placed beacons. Under this formalism, beacons provide an added source of supervision for extracting the task-relevant state (i.e., the features) from robot observations.
@@
Experiments. We conduct multiple simulations and a robot arm experiment in visual imitation learning settings. When the human places the beacon(s) on relevant objects, we find that our RECON algorithm learns the desired task more accurately than prior imitation learning approaches.
@@ IV-C RECON Algorithm
Now that we have defined the model structure (Section IV-A) and loss functions (Section IV-B), we are ready to introduce our RECON algorithm. RECON is an imitation learning approach that extrapolates from the human’s offline dataset. The human teacher first instruments the environment by adding beacons, and then the human provides task demonstrations $\mathcal{D}$. The robot trains its policy network $\pi_{\theta_{1}}$, feature network $f_{\theta_{2}}$, and beacon network $h_{\theta_{3}}$ to minimize the combined loss across the human’s dataset:
@@ V Simulation Experiments
Environments. We evaluate our algorithm in three environments: Static 2D, Robosuite, and Dynamic 2D. The Static 2D environment consists of three objects placed on a line as shown in Figure 3. The robot’s goal is to reach and grasp the object in the center. In each instantiation of the task, the objects are shuffled and positioned randomly. The Robosuite environment has two bins. The left bin contains four objects that are randomly positioned at the start of the task. The robot’s goal is to pick up the right-most object and transfer it to the other bin. In both these environments the objects are static until grasped by the robot. In contrast, the objects in the Dynamic 2D environment move around in a circle (see Figure 4-Left). The robot’s goal is to move away from the red object while disregarding the other objects.
@@ V-A Imitation Learning with Task-Relevant Beacons
Demonstrations. We simulate the robot’s actions for 10 time steps in Static 2D and until task completion in Robosuite, after which the environment is reset. At each step, we record the robot state $x$, the environment state $y$, the beacon reading $b$, and the action $u$. The robot’s state is its 2D position in Static 2D and 6D joint angles in Robosuite. The environment state includes the positions of all objects.
Play data. Since the objects in our environments are static, the beacon readings stay constant throughout the task. As a result, the beacon data recorded with the demonstrations can be insufficient for training the feature and beacon networks. To account for this, we collect an additional dataset of $(y,b)$ pairs by randomizing the environment and placing beacons on the task-relevant objects. Note that we do not collect extra demonstrations. In practice, play data could be collected once to calibrate the beacons and used for teaching multiple tasks.
Independent variables. Our beacons transmit two types of data: distance and position. While distance requires less instrumentation, position is more informative. We separately evaluate our algorithm with both data types: RECON-Distance and RECON-Position. We also evaluate training with and without play data: RECON (Play) and RECON (No Play). As a baseline, we compare RECON to an imitation learning approach that learns a latent representation of the observations [8] and trains a policy conditioned on the robot state and latent values [27]. Similar to RECON, the Baseline includes the policy and feature networks (grey blocks in Figure 2) and trains them by minimizing the loss in Equation (2). However, it does not use any beacon data. We train all methods with the same set of demonstrations.
@@
Results. Our results are summarized in Figure 3. In Static 2D, we test the learned policies in $100$ new task configurations by rolling out the actions for $10$ time steps and measuring the final distance between the robot and the target object. In Robosuite, we measure the mean squared error between the actions of an expert policy and the actions of the learned policies in $110$ task configurations.
@@
In Static 2D, we find that RECON-P achieves a lower final state error than the Baseline, with and without training on the play data. Meanwhile, RECON-D (No Play) performs similar to the Baseline, only surpassing it after training on the play data. RECON-P performs well even without play data because the mapping from the environment state to beacon position, $y\mapsto\phi\mapsto b$, is straightforward — since $y$ already includes the position of the tagged object. As a result, $\phi$ can easily learn to $b$ from just the demonstration data. On the other hand, the mapping from $y$ to distance is indirect and requires additional play data to train $\phi$ effectively.
@@
Figure 4: Results for simulations in the Dynamic 2D environment and the real-world robot experiment. (Left) Simulation results for our experiments in Section V-B averaged over $15$ training and testing runs. We find that robots trained with beacons attached to the task-relevant objects (Exact) achieve the highest rewards while random beacon placement results in the lowest. Though Partial and Other (less relevant) beacon readings also perform worse than Exact, they still outperform the Baseline (which does not use beacons). Exact remains better than the Baseline even with moderate noise ($\sigma=2.5$), only performing worse at high noise levels ($\sigma\geq 6.5$). (Right) Results for our robot experiment in Section VI averaged over $3$ end-to-end runs. With just $4$ demonstrations, a robot trained without beacons (Baseline) fails to recognize the task-relevant objects (blue plate), resulting in poor performance during testing. In contrast, a robot trained with beacon readings (RECON (Play)) learns to focus on the position of the plate and deliver the toast accurately.
@@
In Robosuite, we find that only RECON-P (Play) is able to outperform the Baseline. Unlike Static 2D, where the objects are aligned in a straight line, the objects in this environment are scattered within a bin. Hence, the distance to an object does not always correlate with its position, which explains why the performance does not improve when using beacons that transmit distance instead of position.
@@
Overall these results suggest that when beacon data aligns perfectly with the task-relevant features, our algorithm learns better policies with the same amount of demonstrations compared to learning solely from the environment observations.
@@ V-B Learning from Imperfect Beacon Data
In our previous simulations we assumed that the beacons were accurately placed on the task-relevant objects. However, in real-world settings, users can make mistakes, e.g., placing beacons on irrelevant items. Moreover, these beacon readings can be noisy in practice. How does our algorithm perform with imperfect beacon data?
@@
Demonstrations. We collect demonstrations in the Dynamic 2D environment by randomly initializing the objects and simulating the robot’s actions for 10 time steps. As opposed to the static environments where the observations were vectors of object positions, the observations in this environment are RGB images. Because the objects in this environment move at each time step, we obtain diverse beacon readings in the demonstration data itself. Hence, we do not require any play data for calibrating the beacons for this experiment.
Results. Our results are summarized in Figure 4-Left. We train robot policies using RECON for each beacon placement and noise level. All policies are trained with $10$ demonstrations. We then test the learned policies in $100$ random task configurations by rolling out the actions for $10$ time steps and measuring the final distance between the robot and the red object. Since the robot’s goal is to move away from the red object, the final distance corresponds to its Reward.
As expected, the rewards decrease with increasing noise in the beacon readings. The robot achieves the highest rewards when its policy is trained with perfect beacon data, i.e., exact placement with zero noise. The performance reduces as we use partially or less relevant beacon data, with the lowest rewards for policies trained using randomly placed beacons. Overall, all beacon placements except Random result in higher rewards than the Baseline, which does not use any beacons. These results indicate that the improvement in performance from training with beacons depends on the degree of correlation between all or part of the beacon readings and the task-relevant features.
@@ VI Robot Experiment
In this section we evaluate our algorithm in a real-world environment. Specifically, we train a UR5 robot arm to pick up a piece of toast and drop it on a plate. We compare our RECON approach to the same Baseline as in our simulations.
Experimental setup. The environment includes a blue plate, a red bowl, and a piece of toast placed on a table as shown in Figure 1. The blue plate can lie in two locations: on the left or right edge of the table, while the red bowl is placed anywhere at random. An expert human teacher teleoperates the robot from the toast to the blue plate using a joystick. We collect $2$ examples each for $3$ different plate and bowl configurations — $6$ demonstrations in total. During each demonstration we record the robot joint states $x$, joint velocities $u$, images from a top-down camera $y$, and position readings from a beacon placed near the blue plate $b$. We also collect play data where we attach the beacon to the plate and record $(y,b)$ pairs for various plate and bowl configurations.
Results. Our results are displayed in Figure 4-Right. We train the robot policy with the collected demonstrations and test by executing the learned policy in the real world for $20$ different plate and bowl positions.
After training with only $4$ demonstrations, the Baseline learns to associate the robot’s actions with the positions of both objects. As a result, it fails to generalize when the position of the bowl changes during testing. By contrast, RECON learns to ignore the bowl and focus on the position of the plate by leveraging the beacon readings (especially after training with play data). With more demonstrations, the Baseline gradually learns that the bowl is irrelevant to the task, bringing its performance closer to that of RECON (Play). Overall, these results are consistent with the findings from our simulation experiments and highlight the benefit of incorporating human-placed beacons when teaching robots.
@@ VII Conclusion
In this paper we equipped an expert human teacher with position tracking beacons. The human attached these beacons to relevant objects, and we developed an imitation learning approach that synthesized both the human’s demonstrations and the beacon data to learn the desired task. Specifically, we trained the robot to encode its observations (e.g., camera images) into a compact state representation that maximized mutual information with the beacon data. Our simulations and experiment suggest that this approach mitigates causal confusion, helping the robot to efficiently learn the demonstrated task even when the beacons are removed.

### v2

@@ Abstract
Imitation learning enables robots to learn new tasks from human examples. One fundamental limitation while learning from humans is causal confusion. Causal confusion occurs when the robot’s observations include both task-relevant and extraneous information: for instance, a robot’s camera might see not only the intended goal, but also clutter and changes in lighting within its environment. Because the robot does not know which aspects of its observations are important a priori, it often misinterprets the human’s examples and fails to learn the desired task. To address this issue, we highlight that — while the robot learner may not know what to focus on — the human teacher does. In this paper we propose that the human proactively marks key parts of their task with small, lightweight beacons. Under our framework (RECON) the human attaches these beacons to task-relevant objects before providing demonstrations: as the human shows examples of the task, beacons track the position of marked objects. We then harness this offline beacon data to train a task-relevant state embedding. Specifically, we embed the robot’s observations to a latent state that is correlated with the measured beacon readings: in practice, this causes the robot to autonomously filter out extraneous observations and make decisions based on features learned from the beacon data. Our simulations and a real robot experiment suggest that this framework for human-placed beacons mitigates causal confusion. Indeed, we find that using RECON significantly reduces the number of demonstrations needed to convey the task, lowering the overall time required for human teaching. See videos here: https://youtu.be/oy85xJvtLSU
@@ I Introduction
Leveraging Beacons. We extend imitation learning from offline human demonstrations to include user-placed beacons. Under this formalism, beacons provide an added source of supervision for extracting the task-relevant states (i.e., the features) from robot observations.
@@
Experiments. We conduct multiple simulations and a robot arm experiment in visual imitation learning settings. When the human places the beacon(s) on relevant objects, we find that our RECON algorithm learns the desired task accurately in fewer demonstrations than previous imitation learning approaches. While deploying beacons requires added time, it is considerably less than providing more demonstrations, and thus RECON reduces the total training time.
@@ IV-C RECON Algorithm
Now that we have defined the model structure (Section IV-A) and loss functions (Section IV-B), we are ready to introduce our RECON algorithm. The human teacher first instruments the environment by adding beacons, and then the human provides task demonstrations $\mathcal{D}$. The robot trains its policy network $\pi_{\theta_{1}}$, feature network $f_{\theta_{2}}$, and beacon network $h_{\theta_{3}}$ to minimize the combined loss across the dataset:
@@
We emphasize that our proposed loss in Equation (6) is not constrained to CNNs or MLPs, and designers can replace them with other architectures such as transformers.
@@ V Simulation Experiments
Environments. We evaluate our algorithm in three environments: Static 2D, Robosuite, and Dynamic 2D. The Static 2D environment consists of three objects placed on a line as shown in Figure 3-Left. In each instantiation of the task, the objects are shuffled and positioned randomly. The robot’s goal is to reach and grasp the object in the center. The Robosuite environment has two bins. The left bin contains four objects that are randomly positioned at the start of the task, as shown in Figure 3-Right. The robot’s goal is to pick up the right-most object and transfer it to the other bin. In both Static 2D and Robosuite, the objects are static until grasped by the robot. In contrast, the objects in the Dynamic 2D environment move anticlockwise in a circle (see Figure 4). The robot’s goal is to move away from the red object while disregarding the other objects.
@@ V-A Imitation Learning with Task-Relevant Beacons
Demonstrations. We collect demonstrations by simulating the robot actions until task completion in both environments. At each time step in Static 2D, we record the robot’s position $x$, the positions of the three objects $y$, the beacon reading $b$, and a unit action $u$ in the direction of the central object. Demonstrations in Static 2D have at most 10 time steps. In Robosuite, the actions are generated by an expert policy. At each time step, we record the robot’s 6D joint angles $x$, the position of all objects in the bin $y$, the beacon reading $b$, and joint velocity $u$. After each demonstration, the environments are reset with the objects in random positions.
Play data. Since the objects in our environments are static, the beacon readings stay constant until the robot moves the target object. As a result, the beacon data recorded with the demonstrations can be insufficient for training the feature and beacon networks. To supplement the demonstration data, we collect an additional dataset of $(x,y,b)$ pairs by randomly initializing the environment and recording beacon readings for the task-relevant objects. For example, in Robosuite, we randomize the object positions $y$ and measure the position of the right-most object $b$. We use this play data to further train the feature network with the loss in Equation (5).
Note that we gather play data without simulating the robot, i.e., we do not acquire extra demonstrations. In practice, play data is collected once to learn the mapping $f_{\theta_{2}}(x,y)=\phi$ and used to teach multiple tasks.
Independent variables. As a baseline, we compare RECON to a standard imitation learning approach with feature and policy networks (the grey blocks in Figure 2). The Baseline does not use any beacon data, instead, the feature network is trained end-to-end with the policy network by minimizing the loss Equation (2[27] to learn an unsupervised representation of the environment states.
Our beacons transmit two types of data: distance and position. While distance requires less instrumentation, position is more informative. We separately evaluate our algorithm with both data types: RECON-Distance and RECON-Position. We also evaluate training with and without play data: RECON (Play) and RECON (No Play).
@@
Results. Figure 3 summarizes our results. In Static 2D, we test the learned policies in $100$ random task configurations by rolling out the actions for $10$ time steps and measuring the final distance between the robot and the target object. In Robosuite, we measure the mean squared error between the actions of the expert policy and the actions of the learned policies in $110$ task configurations. The results in both environments are averaged over 20 training runs.
@@
In Static 2D, RECON-P achieves a lower final state error than the Baseline, with and without training on the play data. Meanwhile, RECON-D (No Play) performs similar to the Baseline and requires play data to perform better. RECON-P performs well even without play data because the mapping from the environment state to beacon position, $y\mapsto\phi\mapsto b$, is straightforward — since $y$ already includes the position of the tagged object. As a result, $\phi$ can easily learn to recover $b$ from just the demonstration data. On the other hand, the mapping from $y$ to distance is indirect and requires additional play data to train $\phi$ effectively.
@@
Figure 4: Simulation results averaged over $15$ training and testing runs in the Dynamic 2D environment. We find that robots trained with beacons attached to the task-relevant objects (Exact) achieve the highest rewards while random beacon placement results in the lowest. Though Partial and Other (less relevant) beacon readings also perform worse than Exact, they still outperform the Baseline (which does not use beacons). Exact remains better than the Baseline even with moderate noise ($\sigma=2.5$), only performing worse at high noise levels ($\sigma>4.5$).
@@
In Robosuite, only RECON-P (Play) outperforms the Baseline, while all other approaches have similar action error. Unlike Static 2D, where the objects are aligned in a straight line, the objects in this environment are scattered within a bin. Hence, the robot’s distance to an object does not always correlate with its position, which explains why the performance does not improve when using beacons that transmit distance instead of position.
@@
Overall, these results suggest that when beacon data aligns perfectly with the task-relevant features, our algorithm learns better policies with the same amount of demonstrations compared to learning solely from the environment observations.
@@ V-B Learning from Imperfect Beacon Data
In our previous simulations we assumed that the beacons were accurately placed on the task-relevant objects. However, in real-world settings, users can make mistakes, e.g., placing beacons on irrelevant items. Moreover, these beacon readings can be noisy in practice. We now evaluate how our algorithm performs with imperfect beacon data.
@@
Demonstrations. We collect demonstrations in the Dynamic 2D environment by randomly initializing the objects and simulating the robot’s actions for 10 time steps. As opposed to the static environments where the observations were vectors of object positions, the observations in this environment are RGB images. Because the objects in this environment move at each time step, we obtain diverse beacon readings in the demonstration data itself. Hence, we do not require any play data for training the feature and beacon networks.
Results. Our results are summarized in Figure 4. We train robot policies using RECON for each beacon placement and noise level. All policies are trained with $10$ demonstrations. We then test the learned policies in $100$ random task configurations by rolling out the actions for $10$ time steps and measuring the final distance between the robot and the red object. Since the robot’s goal is to move away from the red object, the final distance corresponds to its Reward.
As expected, the rewards decrease with increasing noise in the beacon readings. The robot achieves the highest rewards when its policy is trained with perfect beacon data, i.e., exact placement with zero noise. The performance reduces as we use less relevant or noisy beacon data, with the lowest rewards for policies trained using randomly placed beacons. Overall, all beacon placements except Random result in higher rewards than the Baseline, which does not use any beacons. These results indicate that the robot’s performance will improve as long as there is some correlation between the beacon readings and the task-relevant features.
@@ VI Robot Experiment
In this section we evaluate our algorithm in a real-world environment. Specifically, we train a UR5 robot arm to pick up a piece of toast and drop it on a plate. We compare our RECON approach to the same Baseline as in our simulations. Unlike our simulated experiments, real-world settings require humans to spend additional time attaching beacons to task-relevant objects. Here we test if deploying physical beacons reduces the overall time required to train the robot.
Figure 5: Results for our robot experiment averaged over $3$ end-to-end runs. With just four demonstrations, a robot trained without beacons (Baseline) fails to recognize the task-relevant objects (blue plate), resulting in poor performance during testing. In contrast, a robot trained with beacon readings (RECON (Play)) learns to focus on the plate and deliver the toast accurately.
Experimental setup. The environment includes a blue plate, a red bowl, and a piece of toast placed on a table as shown in Figure 1. At test time, the blue plate can lie in two locations: on the left or right edge of the table, while the red bowl is placed anywhere at random. An expert human teacher places a beacon near the blue plate, and then teleoperates the robot from the toast to the blue plate using a joystick. During each demonstration we record the robot joint states $x$, joint velocities $u$, images from a top-down camera $y$, and position readings from the beacon near the plate $b$.
Demonstrations. We collect $2$ examples each for $3$ different plate and bowl configurations — $6$ in total. Each demonstration contains ${\sim}150$ $(x,y,b,u)$ samples recorded at 10 Hz. On average it took $16.6$ seconds for a human expert to provide one demonstration and $2.3$ seconds to attach a beacon to the plate. We also collect $4$ seconds of play data by attaching the beacon to the plate and moving the bowl around. The play data contains $40$ $(y,b)$ pairs recorded at $10$ Hz.
Results. Our results are displayed in Figure 5. We train the robot policy with the collected demonstrations and test by executing the learned policy in the real world for $20$ different plate and bowl positions.
After training with only $4$ demonstrations, the Baseline learns to associate the robot’s actions with the positions of both objects. As a result, it fails to generalize when the position of the bowl changes during testing. By contrast, RECON learns to ignore the bowl and focus on the position of the plate by leveraging the beacon readings (especially after training with play data). With more demonstrations, the Baseline gradually learns that the bowl is irrelevant to the task, bringing its performance closer to that of RECON (Play). Overall, these results are consistent with the findings from our simulations and highlight how human-placed beacons can mitigate causal confusion in imitation learning.
Data collection time. To provide $4$ demonstrations, a human working with the Baseline will spend ${\sim}66.4$ seconds interacting with the robot, while with RECON (Play), the human will spend ${\sim}75$ seconds. The extra time includes $4$ seconds of play data and $4.6$ seconds to place the beacon in each of the two positions of the plate. However, without using beacons, the robot does not learn to place the toast accurately after only $4$ demonstrations. The Baseline requires more than $6$ demonstrations to perform similar to RECON (Play), which increases its training time by more than $33.2$ seconds. This shows how RECON reduces the total time required to train robots, even with the added step of attaching beacons.
@@ VII Conclusion
In this paper we equipped an expert human teacher with position tracking beacons. The human attached these beacons to relevant objects, and we developed an imitation learning approach that synthesized both the human’s demonstrations and the beacon data to learn the desired task. Specifically, we trained the robot to encode its observations (e.g., camera images) into a compact state representation that maximized mutual information with the beacon data. Our simulations and experiment suggest that this approach mitigates causal confusion, helping the robot to efficiently learn the demonstrated task and successfully execute it even after the beacons are removed. Future work will explore integration of different beacon types such as auditory and haptic in tasks involving multiple relevant objects and features.